OpenStudy (anonymous):

2(x + 5) > 8x – 8 10 + 8 > 8x – 2x 18> 6x 3> x Did I do this Inequality right or no and if not can u help me and explain what I did wrong and help me get the right answer ?

OpenStudy (anonymous):

you r right hun

OpenStudy (anonymous):

You did it wrong which may result in the wrong outcome (the solution). 2(x+5) > 8x - 8 5*2 = 10 (you got that but you forgot the x)

OpenStudy (anonymous):

@AllTheLonelyKillers She isn't right, please don't give her the wrong impression.

OpenStudy (anonymous):

please explain how u got ur answer

OpenStudy (anonymous):

Her inequality sign needs to be switched. x < 3

OpenStudy (anonymous):

I'm not good with solving inequalities. but she didn't distribute/foil correctly

OpenStudy (anonymous):

what is the answer then

OpenStudy (anonymous):

2(x + 5) > 8x – 8 2x + 10x > 8x-8 12x > 8x-8 Can you take it from here? If not, your step by step might be correct, you just need to flip the sign. SOMEWHERE in inequality step by step, the inequality gets flipped due to the equation. You never flipped it.

OpenStudy (anonymous):

I already said the answer

OpenStudy (anonymous):

well @whalexnuker wouldn't i have to take 12x > 8x - 8 and divide them by 8x or no

OpenStudy (anonymous):

thts exactually what she got x<3

OpenStudy (anonymous):

wait so the only thing I did wrong was not flip the equation

OpenStudy (anonymous):

That's what I'm stuck on XD 12/8 = 1.5 so it doesn't work... honestly, i'd just flip your sign, yes. if the teacher tells you you missed a step, you can always redo it

OpenStudy (anonymous):

x<3=3>x is the same thing

OpenStudy (***[isuru]***):

@whalexnuker U R WRONG!! The answer @NadiaElhaj_1 had got is perfectly correct. the inequality sign will only flipped the side if u multiply or divide the inequality by a negative sign. In this inequality u don't have to do any multiplications or divisions which include negative signs

OpenStudy (anonymous):

thank you good man

OpenStudy (anonymous):

really ? @***[ISURU]*** Thanks

OpenStudy (anonymous):

@***[ISURU]*** @AllTheLonelyKillers I get that it's the same thing but it's correct for a school student to write "x<3" not "3>x"